# Structure growth in $f(Q)$ cosmology

### v1

@@ 5 Results and discussion
In the context of the nonmetric modified theory of gravity, we have performed a thorough statistical analysis resorting to both cosmological expansion history and large-scale structures data. Throughout the investigation, we considered a generic class of power-law models whose gravitational Lagrangian is of the form $f(Q)=Q+\alpha\,Q^{n}$. First, in order to study the cosmic history, in Section 2 MCMC simulations were employed to estimate the best-fit outcomes for the parameters $\Omega_{m}$, $H_{0}$, and $n$. The results were presented in Table 1, drawing upon data from OHD, SNIa, and OHD+SNIa measurements. An extensive statistical analysis, encompassing ${\mathcal{L}(\hat{\theta}|data)}$, $\chi^{2}$, $\chi^{2}$-red(uced), $AIC$, $|\Delta AIC|$, $BIC$, and $|\Delta BIC|$, was carried out (see Table 2). It turned out that for the OHD, SNIa, and OHD+SNIa datasets, the $\Delta AIC$ values providing values of $2.01$, $1.92$, and $0.46$, respectively, indicating a robust support for the proposed $f(Q)$ paradigm. However, the corresponding $\Delta BIC$ values are resulted in $3.05$, $6.87$, and $5.0$, respectively, suggesting, according to Jeffrey’s scale, a comparatively weaker support for the power-law $f(Q)$ gravity model. This prompted us for further statistical investigation to assess the competitiveness of the chosen $f(Q)$ models. The way to follow consisted on utilizing both cosmic expansion measurements and large-scale structure data.
@@
Thus, in Sec. 3 we provided a detailed analysis of the growth of scalar perturbations in $f(Q)$ gravity resorting to the $1+3$ covariant formalism to derive the pertinent equations. As such, the spatial gradients of the evolution equations were presented, and scalar and harmonic decomposition techniques were applied to extract the matter density fluctuations. After having derived the evolution equations, both for the full system without any approximation and their counterparts within the quasi-static approximation, we considered a matter-dominated universe to present the numerical results of the density fluctuations corresponding to both the whole system, i.e., with no simplifying approximation, evolution Eq. (3.26) - shown in the left panel of Fig. 2 - and the quasi-static approximation Eq. (3.29) - shown in the right panel of Fig. 2 - for statistically preferred values of the exponent $n$. Additionally, our analysis allowed us to determine and quantify the discrepancy of theoretical predictions issued from the $\Lambda$CDM and $f(Q)$ power-law model when the density contrast is the variable of interest. To do so, we introduced the dimensionless parameter $\zeta$ as provided in Eq. (3.31). The numerical result of $\zeta(z)$ is also presented in the inner left and right panels of Fig. 2 to see the corresponding amplitude deviations of $\delta(z)$ between $\Lambda$CDM and $f(Q)$ power-law model by considering the solution of the full system and quasi-static approximations. We were also capable of comparing the density fluctuations predictions when obtained from either the quasi-static approximation or the full system, as represented in Fig. 3. Such a figure shows that in the context of f(Q) theories, the quasi-static approximation provides a faithful representation of the full resolution, with consistency levels of 84.89% at $n=0.026$, 82.1% at $n=0.09$, and 76.98% at $n=0.134$ in the $f(Q)$ gravity power-law model.
@@
Acknowledgements
The authors would like to thank Javier de Cruz Pérez, Renier Hough, and Antonio L. Maroto for insightful comments during the preparation of the manuscript. Endalkachew Tsegaye was initially involved in earlier aspects of the work, and his contribution is duly acknowledged. AA acknowledges that this work was part of the research programme “New Insights into Astrophysics and Cosmology with Theoretical Models Confronting Observational Data” of the National Institute for Theoretical and Computational Sciences of South Africa. AdlCD acknowledges support from BG20/00236 action (MCINU, Spain), NRF Grant CSUR23042798041, CSIC Grant COOPB23096, and Grant PID2021-122938NB-I00 funded by MCIN/AEI/10.13039/501100011033 and by ERDF A way of making Europe.
@@ Appendix A Growth structure data
SDSS 0.10 0.464 $\pm$ 0.040 0.769 $\pm$ 0.105 [40]
Vipers PDR-2 0.60 0.93 $\pm$ 0.22 0.52 $\pm$ 0.06 [30, 41]
Vipers PDR-2 0.86 0.99 $\pm$ 0.19 0.48 $\pm$ 0.04 [30, 41]
@@
2MTF 0.001 0.505 $\pm$ 0.085 [42]
6dFGS+SNIa 0.02 0.428 $\pm$ 0.0465 [43]
IRAS+SNIa 0.02 0.398 $\pm$ 0.065 [44, 45]
2MASS 0.02 0.314 $\pm$ 0.048 [44, 46]
SDSS 0.10 0.376 $\pm$ 0.038 [40]
SDSS-MGS 0.15 0.490$\pm$ 0.145 [47]
2dFGRS 0.17 0.510 $\pm$ 0.060 [48]
GAMA 0.18 0.360 $\pm$ 0.090 [49]
GAMA 0.38 0.440 $\pm$ 0.060 [49]
SDSS-LRG-200 0.25 0.3512 $\pm$ 0.0583 [50]
SDSS-LRG-200 0.37 0.4602 $\pm$ 0.0378 [50]
BOSS DR12 0.31 0.469 $\pm$ 0.098 [51]
BOSS DR12 0.36 0.474 $\pm$ 0.097 [51]
BOSS DR12 0.40 0.473 $\pm$ 0.086 [51]
BOSS DR12 0.44 0.481 $\pm$ 0.076 [51]
BOSS DR12 0.48 0.482 $\pm$ 0.067 [51]
BOSS DR12 0.52 0.488 $\pm$ 0.065 [51]
BOSS DR12 0.56 0.482 $\pm$ 0.067 [51]
BOSS DR12 0.59 0.481 $\pm$ 0.066 [51]
BOSS DR12 0.64 0.486 $\pm$ 0.070 [51]
WiggleZ 0.44 0.413 $\pm$ 0.080 [52]
WiggleZ 0.60 0.390 $\pm$ 0.063 [52]
WiggleZ 0.73 0.437 $\pm$ 0.072 [52]
Vipers PDR-2 0.60 0.550 $\pm$ 0.120 [30, 41]
Vipers PDR-2 0.86 0.400 $\pm$ 0.110 [30, 41]
FastSound 1.40 0.482 $\pm$ 0.116 [53]
SDSS-IV 0.978 0.379 $\pm$ 0.176 [54]
SDSS-IV 1.23 0.385 $\pm$ 0.099 [54]
SDSS-IV 1.526 0.342 $\pm$ 0.070 [54]
SDSS-IV 1.944 0.364 $\pm$ 0.106 [54]

### v4

@@ 5 Conclusions
First, in order to study the cosmic history, in Section 2 MCMC simulations were employed to estimate the best-fit outcomes for the parameters $\Omega_{m}$, $H_{0}$, and $n$. The results were presented in Table 1, drawing upon data from OHD, SNIa, and OHD+SNIa measurements. Therein OHD data provide their best fit for exponent $n>0$, whereas both SNIa and SNIa+OHD provide a similar negative value of $n$ as their best fit. Then, an extensive statistical analysis, encompassing ${\mathcal{L}(\hat{\theta}|data)}$, $\chi^{2}$, $\chi^{2}$-red(uced), $\rm{AIC}$, $\Delta\rm{AIC}$, $\rm{BIC}$, and $\Delta\rm{BIC}$, was carried out (see Table 2). It turned out that for the OHD, SNIa, and OHD+SNIa datasets, the $\Delta\rm{AIC}$ estimator provided values of $1.46$, $1.74$, and $1.42$, respectively, indicating robust support for the proposed $f(Q)$ paradigm. However, the $\Delta\rm{BIC}$ values corresponded to $3.21$, $7.20$, and $5.88$, respectively, suggesting comparatively weaker support for the power-law $f(Q)$ gravity model. This apparent contradiction prompted us to assess the competitiveness of the chosen $f(Q)$ models resorting to further statistical investigation. The way to follow consisted of utilizing both cosmic expansion measurements and large-scale structure data.
@@
Next, in Section 3 we provided a detailed analysis of the scalar perturbations equations in $f(Q)$ gravity resorting to the $1+3$ covariant formalism. As such, the spatial gradients of the evolution equations were presented, and scalar and harmonic decomposition techniques were applied to extract the matter density fluctuations. After having derived the evolution equations, both for the full system without any approximation and their counterparts within the quasi-static approximation, we considered a matter-dominated universe to present the numerical results of the density fluctuations corresponding to both the whole system, i.e., with no simplifying approximation, governed by Eq. (41) - shown in the left panel of Fig. 2 - and the quasi-static approximation by means of Eq. (3.1) - shown in the right panel of Fig. 2 - for statistically preferred values of the exponent $n$. Additionally, our analysis allowed us to both determine and quantify the discrepancy of theoretical predictions issued from the $\Lambda$CDM and $f(Q)$ power-law model when the density contrast is the variable of interest. To do so, we introduced the dimensionless parameter $\zeta$ as provided in Eq. (46). The numerical results of $\zeta(z)$ were also presented in the inner left and right panels of Fig. 2 to see the corresponding amplitude deviations of $\delta(z)$ between $\Lambda$CDM and $f(Q)$ power-law model by considering either the solution of the full system or that within the quasi-static approximation. We were also capable of comparing the density fluctuations predictions when obtained from either the quasi-static approximation or the full system, as represented in Fig. 3. Such a figure served us to claim that in the context of $f(Q)$ gravity power-law models, the quasi-static approximation provides a faithful representation of the full resolution, with consistency levels of 94.03% at $n=0.215$, 98.7% at $n=-0.085$, and 96.61% at $n=-0.275$.
@@
Acknowledgments
The authors would like to thank Javier de Cruz Pérez, Renier Hough, and Antonio L. Maroto for insightful comments during the preparation of the manuscript. Endalkachew Tsegaye was initially involved in earlier aspects of the work, and his contribution is duly acknowledged. AA acknowledges that this work was part of the research programme “New Insights into Astrophysics and Cosmology with Theoretical Models Confronting Observational Data” of the National Institute for Theoretical and Computational Sciences of South Africa. AdlCD acknowledges support from BG20/00236 action (MCINU, Spain), NRF Grant CSUR23042798041, CSIC Grant COOPB23096, Project SA097P24 funded by Junta de Castilla y León (Spain) and Grant PID2021-122938NB-I00 funded by MCIN/AEI/10.13039/501100011033 and by ERDF A way of making Europe.
Data Availability
There are no associated data with this article. No new data were generated or analysed in support of this research.
@@ Appendix A Growth structure data
SDSS 0.10 0.464 $\pm$ 0.040 0.769 $\pm$ 0.105 Shi (2019)
Vipers PDR-2 0.60 0.93 $\pm$ 0.22 0.52 $\pm$ 0.06 De La Torre et al. (2017); Pezzotta et al. (2017)
Vipers PDR-2 0.86 0.99 $\pm$ 0.19 0.48 $\pm$ 0.04 De La Torre et al. (2017); Pezzotta et al. (2017)
@@
2MTF 0.001 0.505 $\pm$ 0.085 Howlett et al. (2017)
6dFGS+SNIa 0.02 0.428 $\pm$ 0.0465 Huterer et al. (2017)
IRAS+SNIa 0.02 0.398 $\pm$ 0.065 Hudson & Turnbull (2012); Turnbull et al. (2012)
2MASS 0.02 0.314 $\pm$ 0.048 Hudson & Turnbull (2012); Davis et al. (2011)
SDSS 0.10 0.376 $\pm$ 0.038 Shi (2019)
SDSS-MGS 0.15 0.490$\pm$ 0.145 Howlett et al. (2015)
2dFGRS 0.17 0.510 $\pm$ 0.060 Song & Percival (2009)
GAMA 0.18 0.360 $\pm$ 0.090 Blake et al. (2013)
GAMA 0.38 0.440 $\pm$ 0.060 Blake et al. (2013)
SDSS-LRG-200 0.25 0.3512 $\pm$ 0.0583 Samushia et al. (2012)
SDSS-LRG-200 0.37 0.4602 $\pm$ 0.0378 Samushia et al. (2012)
BOSS DR12 0.31 0.469 $\pm$ 0.098 Wang et al. (2018b)
BOSS DR12 0.36 0.474 $\pm$ 0.097 Wang et al. (2018b)
BOSS DR12 0.40 0.473 $\pm$ 0.086 Wang et al. (2018b)
BOSS DR12 0.44 0.481 $\pm$ 0.076 Wang et al. (2018b)
BOSS DR12 0.48 0.482 $\pm$ 0.067 Wang et al. (2018b)
BOSS DR12 0.52 0.488 $\pm$ 0.065 Wang et al. (2018b)
BOSS DR12 0.56 0.482 $\pm$ 0.067 Wang et al. (2018b)
BOSS DR12 0.59 0.481 $\pm$ 0.066 Wang et al. (2018b)
BOSS DR12 0.64 0.486 $\pm$ 0.070 Wang et al. (2018b)
WiggleZ 0.44 0.413 $\pm$ 0.080 Blake et al. (2012)
WiggleZ 0.60 0.390 $\pm$ 0.063 Blake et al. (2012)
WiggleZ 0.73 0.437 $\pm$ 0.072 Blake et al. (2012)
Vipers PDR-2 0.60 0.550 $\pm$ 0.120 De La Torre et al. (2017); Pezzotta et al. (2017)
Vipers PDR-2 0.86 0.400 $\pm$ 0.110 De La Torre et al. (2017); Pezzotta et al. (2017)
FastSound 1.40 0.482 $\pm$ 0.116 Okumura et al. (2016)
SDSS-IV 0.978 0.379 $\pm$ 0.176 Wang et al. (2018a)
SDSS-IV 1.23 0.385 $\pm$ 0.099 Wang et al. (2018a)
SDSS-IV 1.526 0.342 $\pm$ 0.070 Wang et al. (2018a)
SDSS-IV 1.944 0.364 $\pm$ 0.106 Wang et al. (2018a)